OpenStudy (anonymous):

In a word game, you choose a tile from a bag, replace it, and then choose another. If there are 14 vowels and 32 consonants, what is the probability you will choose a consonant then a vowel ?

OpenStudy (anonymous):

a) 224/1,035 b) 112/529 c) 529/112 d) 9/23

OpenStudy (anonymous):

@robz8

OpenStudy (anonymous):

14 + 32 = 46 there are 46 possible choices if the odds of pulling a vowel are 14/46 and a consonant 32/46 then the odds of pulling a consonant then a vowel are 32/46 * 14/46 \[\frac{ 32 }{ 46 } * \frac{14}{46}= \frac{ 448}{ 2116}\]

OpenStudy (anonymous):

we can reduce this though \[\frac{ 448 }{ 2116 } = \frac{ 112 }{ 529 }\]

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

you're welcome